# On the existence, uniqueness and stability of solutions of SDEs with state-dependent variable exponent

### v3

@@ 2 Preliminaries and auxiliary results
Remark 1.
The function $h(x)=1+\frac{0.5}{1+x}$ satisfies the hypotheses $(\mathbf{h_{1}})$-$(\mathbf{h_{3}})$. $(\mathbf{h_{1}})$ and $(\mathbf{h_{2}})$ easily follows. If one lets $\delta=1$, $M_{0}\geq 0.5$, $C_{0}\geq 1$ and $\alpha\geq 1$ then $(\mathbf{h_{3}})$ is also satisfied.
@@
Using the elementary calculus along with the assumptions $(\mathbf{h_{1}})$ and $(\mathbf{h_{3}})$, it follows
@@ 4.2 The asymptotic estimate of the solution
$\limsup_{t\to\infty}\frac{1}{t}\log(X(t))\leq\hat{K}$, (37)
@@
$\limsup_{n\to\infty}\frac{1}{t}\log(X(t))$ $\leq\limsup_{n\to\infty}\frac{1}{2t}\log(1+X^{2}(t))$
@@ 4.3 The stability of the solution
Applying the Burkholder-Davis-Gundy (BDG) inequality and the Hölder inequality together yields
@@ 5 Application
SDEs have numerous applications, one of the most notable being their role in representing solutions to partial differential equations (PDEs) through the Feynman–Kac formula. This formula establishes a fundamental connection between SDEs and PDEs, providing a probabilistic framework for analyzing and solving partial differential equations.
@@
where $f$ is a real-valued $C^{1}$ function; $0<\mu,\sigma$ and $c\geq 0$ real parameters; $L$ is the linear differential operator
@@
Suppose $u$ is the unique solution to Poisson’s equation (56) in $\mathcal{O}$. Then $u$ is given by Feynman–Kac formula

### v4

@@ 2 Preliminaries and auxiliary results
2.1 Illustrative examples
We present several examples of state-dependent exponent functions that satisfy assumptions $(\mathbf{h_{1}})-(\mathbf{h_{3}})$ and therefore fall within the framework of the model ($\mathcal{GM}$}). These examples illustrate how the proposed formulation extends classical SDE models by allowing the elasticity of the drift and diffusion terms to vary with the state variable.
Example 1
Let
$p(x)=1+\frac{a}{1+x},\qquad q(x)=1+\frac{b}{1+x},\qquad a,b>0.$
Then ($\mathcal{GM}$}) becomes
$dX(t)=\mu(t)X(t)^{1+\frac{a}{1+X(t)}}dt+\sigma(t)X(t)^{1+\frac{b}{1+X(t)}}dW(t).$
Since $p(x),q(x)\to 1$ as $x\to\infty$, the dynamics asymptotically approach the GBM, while for small values of $x$ the model exhibits stronger nonlinear drift and diffusion effects. Such a structure can be useful for modeling systems where growth and variability are amplified when the state variable is small.
Example 2
Let
$p(x)=1,\qquad q(x)=\gamma+\frac{c}{1+x},$
where $\gamma>0$ and $c>0$. Then ($\mathcal{GM}$}) becomes
$dX(t)=\mu(t)X(t)\,dt+\sigma(t)X(t)^{\gamma+\frac{c}{1+X(t)}}dW(t).$
In this case the diffusion elasticity varies with the state variable but approaches the constant value $\gamma$ as $x\to\infty$. Hence the model can be viewed as a state-dependent generalization of the classical CEV process.
Example 3
Let
$p(x)=1+\frac{a}{1+x^{2}},\qquad q(x)=1+\frac{b}{1+x^{2}},\qquad a,b>0.$
Then ($\mathcal{GM}$}) becomes
$dX(t)=\mu(t)X(t)^{1+\frac{a}{1+X(t)^{2}}}dt+\sigma(t)X(t)^{1+\frac{b}{1+X(t)^{2}}}dW(t).$
Here the nonlinear effects are most pronounced near the origin and gradually vanish as the state grows. Consequently the model behaves like a GBM for large values of $x$ while allowing richer nonlinear dynamics for moderate and small values of the state variable.
Next, we will continue with some basic concepts of the measure-theoretic probability and the theory of stochastic process.
@@ 2.1 Illustrative examples
Using the assumptions $(\mathbf{h_{1}})$ and $(\mathbf{h_{3}})$, it follows
@@ 4.2 The asymptotic estimate of the solution
$\limsup_{t\to\infty}\frac{\log(X(t))}{t}\leq\hat{K}$, (37)
@@
[…]
@@ 4.3 The stability of the solution
Applying the Burkholder-Davis-Gundy inequality [5, 7] and the Hölder inequality together yields
@@ 5 Application
SDEs have numerous applications, one of the most notable being their role in representing solutions to (deterministic) differential equations through the Feynman–Kac formula. This formula provides a fundamental link between SDEs and differential equations, offering a probabilistic framework for analyzing and solving such equations.
@@
where $f$ is a real-valued $C^{1}$ function; $\mu,\sigma>0$ and $c\geq 0$ real parameters; $L$ is the linear differential operator
@@
Suppose $u$ is the unique solution to Poisson’s equation (56) in $\mathcal{O}$. Then $u$ is given by the Feynman–Kac formula